From a regular deck off 52 cards, what is the probability of choosing a card that isn't an ace or a face card, but that is divisible by 2?
Answer by Theo(13342)

a regular deck of cards has the following:

4 suits (spades, clubs, diamonds, hearts)

4 kings (one of each suit)
4 queens (one of each suit)
4 jacks (one of each suit)
4 aces (one of each suit)
36 cards numbered 2 through 10 (9 of each suit)

if the numbered cards are divisible by 2, then they are even.

that would be 2,4,6,8,10 = 5 of each suit * 4 suits = 20 total.

the probability of getting a card that isn't a face card and isn't an ace and is divisible by 2 is therefore 20/52 = 10/26 = 5/13.